# Factuality Beyond Coherence: Evaluating LLM Watermarking Methods for Medical Texts

### v1

@@ Abstract
As large language models (LLMs) adapted to sensitive domains such as medicine, their fluency raises safety risks, particularly regarding provenance and accountability. Watermarking embeds detectable patterns to mitigate these risks, yet its reliability in medical contexts remains untested. Existing benchmarks focus on detection-quality tradeoffs, overlooking factual risks under low-entropy settings often exploited by watermarking’s reweighting strategy.
@@
1 Introduction
LLMs have advanced human-like text generation capability, raising concerns about potentially harmful or biased information in various use case, including in the medical domain [xue2023trojllm, ghanim-etal-2024-jailbreaking, xue-etal-2024-badfair, zhang2025towards]. Watermarking techniques serve as a safeguard by embedding subtle statistical patterns into generated content that enable detection while preserving quality [kirchenbauer2023watermark, gu2023learnability, kuditipudi2024robust].
While watermarking methods show promise for securing information in LLM outputs, their effectiveness in the medical domain is still underexplored [kong2024protecting]. Moreover, watermarking techniques suffer from issues such as lack of awareness of critical low-entropy tokens [lee-etal-2024-wrote], which could distort medical terminology and compromise factual integrity.
Current watermarking evaluation limitations primarily asses detection-quality tradeoffs from a domain-agnostic perspective, where quality assessment using human or LLM evaluators focuses on measuring text preference [tu-etal-2024-waterbench, molenda-etal-2024-waterjudge] or semantic coherence [singh2024new], remaining unaware of domain-factuality details.
Our analysis reveals critical trade-offs in watermarking methods for medical text. We find that generation-time watermarks achieve high detection rates but alter token entropy distributions, particularly affecting medical entities. Our evaluation workflow demonstrates that factuality degradation is often more severe than coherence loss, a distinction missed by standard quality metrics. Human evaluations show stronger agreement with our FWS. These results highlight the need for tailored watermarking approaches in medical domains where factual integrity is crucial.
@@ 1 Introduction
[…]
@@ 2 Related Work
Watermarking methods for LLMs.  Methods to detect LLM-generated text has been studied in previous works, like leveraging the statistical properties without changing the text itself [mitchell2023detectgpt, gehrmann2019gltr], adding lexical feature as a pattern to distinguish [he2022protecting], as well as training both traditional and neural-based classifier to classify human-generated and model-generated texts [openai, guo2023close, singh2024new]. Recently, a series of work embeds watermarks within LLM-generated texts by either modifying logits from the LLM [kirchenbauer2023watermark, kirchenbauer2024reliability], adding a selective technique tailored for domain requirements [lee-etal-2024-wrote], manipulating the sampling procedure [christ2024undetectable, kuditipudi2024robust], or investigate the learnability of watermarks in the distillation process [gu2023learnability].
@@
Benchmarking watermarking methods.  Existing benchmarking for watermarking methods has defined evaluation criteria, particularly in general-domain scenarios. Most benchmarks emphasize detectability to ensure watermark robustness [tu-etal-2024-waterbench]and complement this with quality assessments often based on LLM-as-evaluator preferences [singh2024new]. However, these evaluations typically lack domain-specific considerations and rely heavily on automatic metrics [ajith2024downstream] or simplified pairwise comparisons, with no further human validation or detailed justification of results [molenda-etal-2024-waterjudge].
@@
Securing medical LLM’s. The development of medical language models supports the applicability of various tasks in broader clinical and healthcare settings. Despite this progress, efforts to incorporate safeguards for medical language models remain limited, such as in the jailbreaking case or model watermarking via backdoor [zhang2025towards, kong2024protecting]. Given the relatively recent development of LLM watermarking techniques, their applicability and effectiveness in critical domains like medical remain unexplored and need further investigation.
3 Evaluation Workflow
General workflow  Our framework evaluates watermarking methods across three key quality aspects, illustrated in Figure 1. 1) Coherence captures the fluency of the generated text and is commonly used in watermarking evaluation. 2) Relevance/Completeness measures whether the context surrounding critical medical terminology is preserved, ensuring watermarking mechanisms do not distort model confidence. [^1] Variations in symptom descriptions may correspond to distinct diagnoses, a model should not fluctuate in confidence when generating crucial term like ”cancer”. See Figure  4. 3) Factual accuracy assesses whether newly introduced medical terminology preserves correct semantics to avoid hallucination  [^2] Entity errors source to factuality hallucination [li-etal-2024-dawn]. See Table 7 for rate and Table 13 for example.. The latter two aspects emphasize potential factual corruption, with detailed analysis provided in Section 5.4.
These aspects are implemented and applicable through two approaches:
  • Traditional automatic metrics. Our framework also integrates established metrics to enable unified quality quantification. Coherence is measured using non-reference metrics such as Perplexity [kirchenbauer2023watermark] or reference-based similarity measures such as SimCSE [huo2024token]. Accuracy is evaluated with task-oriented, reference-based metrics including ROUGE-2, ROUGE-L, F1 [he2024can], and factuality metric AlignScore [zha2023alignscore], using natural text as the ground truth.
Factuality-weighted Score (FWS)  The propose aspects allow us to quantify the overall quality metric as shown in Equation 1. To prioritize relevant context and factual accuracy over coherence, we set $\alpha=0.4$ and $\beta=0.2$ to indicate that Relevance or Completeness ($Rel.$) and Accuracy ($FactAcc.$) are equally important, and they are twice as important as Coherence ($Coh.$). This weighting was based on empirical consideration through sensitivity analysis provided in Table 5.
$FWS=\alpha(\text{Rel}+\text{FactAcc})+\beta(\text{Coh})$ (1)
Human evaluation
To evaluate both automatic metrics and GPT-Judger ratings on equal footing, we conducted a human evaluation with 6 respondents, including graduate students and medical practitioners, ensuring familiarity with LLMs and medical expertise. The evaluation focused on the KGW watermarking method applied to the Question Answering (QA) task, chosen as KGW underlies many LLM watermarking approaches and QA represents the most realistic practical use case for medical language models. Each item was rated by 3 respondents to obtain an average score using the quality aspects in Table 1. The full evaluation template is provided in Appendix B.2.
Pearson correlation was computed to measure agreement between human judgments and both traditional automatic metrics and GPT-Judger ratings. Prior to this, a Nemenyi test was applied to detect significant differences among the proposed aspects, with lower p-values indicating greater distinction [fu2024qgeval].
@@ 4.2 Watermarking Methods
We evaluate six watermarking methods across two categories [liu2024survey]. Post-hoc watermarking applies after text generation. We evaluate two schemes from this category: LogRank [gehrmann2019gltr], which classifies text by thresholding the average log-rank of tokens, as machine-generated text typically has smaller average log-ranks, and DetectGPT [mitchell2023detectgpt], which exploits the tendency of model-generated text to occupy negative curvature regions in the log probability function by comparing original passages with semantically similar perturbations.
The second category is generation-time watermarking, where watermarks are embedded during the LLM’s text generation process. We evaluate four approaches: KGW [kirchenbauer2023watermark], a foundational logit-based method that partitions the vocabulary into red and green lists based on a previous token hash; SWEET [lee-etal-2024-wrote], an improvement of the KGW method on addressing low entropy tokens for code generation domain However, another work [huo2024token] showed that SWEET works well in the general domain, thus we believe potentially useful in medical text where low entropy tokens can be found; DiPmark [wu2024resilient] a distribution-preserving watermark that uses complementary reweighting strategies with random token permutations to maintain original text quality; and EXP-edit [kuditipudi2024robust], a token sampling-based method modifying probabilities based on token rank.
This selection allows comparison across three strategies post-hoc (PH), logit-based (LB), distribution preserving or distortion-free (DF), while also includes methods addressing potential domain-specific challenges like low entropy. All generation-time watermarks used the MarkLLM toolkit [pan-etal-2024-markllm] with default parameters. Brief methods are available in Appendix B.4, and parameter details are available in Appendix A.2, Table 10.
4.3 Tasks and Datasets
To assess watermark performance in realistic medical scenarios, we employ three distinct tasks using established medical datasets. First, Text Completion using the HealthQA dataset [zhu2019hierarchical] requires models to complete medical passages (approx. 200 tokens), evaluating the impact on fundamental coherence and fluency. Second, Question Answering (QA), also on HealthQA, challenges models to answer medical questions (approx. 200 tokens), specifically testing factual accuracy preservation under watermarking. Third, Summarization using the MeQSum dataset [abacha2019summarization] involves generating concise summaries (< 100 tokens) of medical questions, evaluating performance on challenging short-text generation [kirchenbauer2023watermark] and the crucial preservation of essential factual information. These tasks represent diverse medical tasks and output lengths, offering a comprehensive view of watermark impact on text quality and factual integrity in the medical domain. Details on how we built each task are available in Appendix B.3.
4.4 Model and Hardware
We use Meditron-7B [chen2023meditron70b] [^3] https://huggingface.co/epfl-llm/meditron-7b as the main model for most experiments. This model included training on clinical guidelines, more suitable for consumer QA, unlike most models, which limited their training on abstracts and articles. Meditron is a medical adaptation model from Llama-2-7B [touvron2023llama] through continued pretraining on the medical GAP-Replay corpus. We also evaluate other medical models like MedLlama-3-8B [^4] https://huggingface.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0 and BioMistral 7B [^5] https://huggingface.co/BioMistral/BioMistral-7B [labrak2024biomistral]. Though this variation is not our main focus, we only include RQ1 results using these models and provide additional analysis in Appendix A.1. All experiments are conducted on a workstation equipped with an AMD Ryzen Threadripper PRO 3955WX (16-Cores) processor and two NVIDIA GeForce RTX 3090 GPUs (24GB VRAM each).
5 Results
@@ 5.1 Detectability and Quality (RQ1)
Table 2 presents comparative performance of watermarking methods across generation tasks, measuring both detection capability (TPR, AUROC) and text quality preservation (PPL, SimCSE). The results show a clear trade-off between detectability and output quality among different watermarking approaches.
@@ 5.2 Task-oriented Performance (RQ2)
Table 3 presents reference-based quality metrics evaluating generated text against natural text ground-truth, deliberately to give insights of factual preservation across different tasks. Unlike the general quality metrics in RQ1, task-specific metrics (ROUGE-2, ROUGE-L, F1, and AlignScore) assess how watermarking affects bigram preservation, subsequence matching, and precision-recall balance.
@@ 5.3 Factuality-weighted Score FWS (RQ3)
For the task-specific performance across watermarking approaches: SWEET demonstrates superior performance in Text Completion (0.556), while KGW achieves the highest scores in both Question Answering (0.411) and Summarization (0.419). EXP-edit consistently shows the weakest performance across all tasks. This is likely due to its alternative token sampling strategy, which causes the model to behave almost like a different language model compared to the un-watermarked version. In this case, it introduces more drawbacks than the intended benefit of distortion-free watermark, as claimed [kuditipudi2024robust].
@@
Human Evaluation  Prior step of evaluating human ratings done by conducting Nemenyi test and calculating p-value significance on quality aspects define in Table 1. P-value results for the following pairs (Coherence, Fact_1), (Coherence, Fact_2), and (Fact_1, Fact_2) are 0.635, 0.001, and 0.004 respectively. The proposed quality aspects exhibit distinct characteristics consistent with intuition, underscoring the role of detailed factuality aspects in the medical domain. Furthermore, our sensitivity analysis of Equation 1 shows consistent trend where the more FWS emphasizes factuality, the better it correlates to human judgment. While the 2:1 weighting did not yield the strongest correlation, it outperformed coherence-heavy settings and was chosen as an intuitive benchmark for a factuality-heavy configuration as seen in Table 5.
Configuration FWS Param. ($\alpha$, $\beta$) Auto- Human GPTJ- Human
Coherence-Heavy (1:2) (0.25, 0.5) 0.089 0.814
Equal Weighting (1:1) (0.33, 0.33) 0.181 0.833
Current (Ours) (2:1) (0.4, 0.2) 0.248 0.840
Factuality-Heavy (4:1) (0.44, 0.11) 0.276 0.841
Factuality-Heavy (6:1) (0.46, 0.08) 0.283 0.841

### v2

@@ Abstract
As large language models (LLMs) are adapted to sensitive domains such as medicine, their fluency raises safety risks, particularly regarding provenance and accountability. Watermarking embeds detectable patterns to mitigate these risks, yet its reliability in medical contexts remains untested. Existing benchmarks focus on detection-quality tradeoffs and overlook factual risks. In medical text, watermarking often reweights low-entropy tokens, which are highly predictable and often carry critical medical terminology. Shifting these tokens can cause inaccuracy and hallucinations, risks that prior general-domain benchmarks fail to capture.
@@
Factuality Beyond Coherence: Evaluating LLM Watermarking Methods for Medical Texts
Rochana Prih Hastuti, Rian Adam Rajagede, Mansour Al Ghanim, Mengxin Zheng, Qian Lou University of Central Florida Orlando, FL {rochana, rian, mansour.alghanim, mengxin.zheng, qian.lou}@ucf.edu
1 Introduction
LLMs have advanced human-like text generation capability, raising concerns about potentially harmful or biased information in various use case, including in the medical domain Xue et al. (2024); Zhang et al. (2025); Al Ghanim et al. (2023); Lou et al. ; Hastuti et al. (2025). Watermarking techniques serve as a safeguard by embedding subtle statistical patterns into generated content that enable detection while preserving quality Kirchenbauer et al. (2023); Gu et al. (2023); Kuditipudi et al. (2024).
While watermarking methods show promise for securing information in LLM outputs, their effectiveness in the medical domain remains underexplored Kong et al. (2024). A key limitation is that watermarking often treats all tokens uniformly, ignoring low-entropy tokens Lee et al. (2024). These are highly predictable words that often carry critical factual content in medical text, such as disease names. Reweighting them to embed a watermark can alter their meaning, compromising factual accuracy and increasing the risk of hallucinations. Current watermarking evaluation primarily assesses detection-quality tradeoffs from a domain-agnostic perspective, where quality assessment using human or LLM evaluators focuses on measuring text preference Tu et al. (2024); Molenda et al. (2024) or semantic coherence Singh and Zou (2024), without considering the domain-specific factual risks associated with low-entropy tokens
Our analysis reveals critical trade-offs in watermarking methods for medical text. We find that generation-time watermarks, where watermarks are embedded during the LLM’s text generation process, achieve high detection rates but alter token entropy distributions, particularly affecting medical entities. Our evaluation workflow demonstrates that factuality degradation is often more severe than coherence loss, a distinction missed by standard quality metrics. Human evaluations show stronger agreement with our FWS. These results highlight the need for tailored watermarking approaches in medical domains where factual integrity is crucial.
@@ 1 Introduction
The code and experimental data are available at: https://github.com/rochanaph/fact-eval-wllm
@@ 2 Related Work
Watermarking methods for LLMs.  Methods to detect LLM-generated text has been studied in previous works, like leveraging the statistical properties without changing the text itself Mitchell et al. (2023); Gehrmann et al. (2019), adding lexical feature as a pattern to distinguish He et al. (2022), as well as training both traditional and neural-based classifier to classify human-generated and model-generated texts OpenAI (2023); Guo et al. (2023); Singh and Zou (2024). Recently, a series of work embeds watermarks within LLM-generated texts by either modifying logits from the LLM Kirchenbauer et al. (2023, 2024), adding a selective technique tailored for domain requirements Lee et al. (2024), manipulating the sampling procedure Christ et al. (2024); Kuditipudi et al. (2024), or investigate the learnability of watermarks in the distillation process Gu et al. (2023).
@@
Benchmarking watermarking methods.  Existing benchmarking for watermarking methods has defined evaluation criteria, particularly in general-domain scenarios. Most benchmarks emphasize detectability to ensure watermark robustness Tu et al. (2024) and complement this with quality assessments often based on LLM-as-evaluator preferences Singh and Zou (2024). However, these evaluations typically lack domain-specific considerations and rely heavily on automatic metrics Ajith et al. (2024) or simplified pairwise comparisons, with no further human validation or detailed justification of results Molenda et al. (2024).
Securing medical LLM’s. The development of medical language models supports the applicability of various tasks in broader clinical and healthcare settings. Despite this progress, efforts to incorporate safeguards for medical language models remain limited, such as in the jailbreaking case or model watermarking via backdoor Kong et al. (2024); Xue et al. (2023); Al Ghanim et al. (2024); Hastuti et al. (2025); Zheng et al. (2024). Given the relatively recent development of LLM watermarking techniques, their applicability and effectiveness in critical domains like medical remain unexplored and need further investigation.
3 Evaluation Workflow
General workflow  Our framework is illustrated in Figure 1. First, we feed the Medical LLMs with a prompt such as a medical question, and then we evaluate the quality of its watermarked and unwatermarked output. We evaluate watermarking methods across three key quality aspects. 1) Coherence captures the fluency of the generated text and is commonly used in watermarking evaluation. 2) Relevance/Completeness measures whether the context surrounding critical medical terminology is preserved, ensuring watermarking mechanisms do not distort model confidence. [^1] Variations in symptom descriptions may correspond to distinct diagnoses, a model should not fluctuate in confidence when generating crucial terms like ”cancer”. See Figure  4. 3) Factual accuracy assesses whether newly introduced medical terminology preserves correct semantics to avoid hallucination  [^2] Entity errors source to factuality hallucination Li et al. (2024). See Table 7 for rate and Table 13 for example.. The latter two aspects emphasize potential factual corruption, with detailed analysis provided in Section 5.4.
In Step 1, we implement and evaluate these three quality aspects through two approaches:
  • Traditional automatic metrics. Our framework also integrates established metrics to enable unified quality quantification. Coherence is measured using non-reference metrics such as Perplexity Kirchenbauer et al. (2023) or reference-based similarity measures such as SimCSE Huo et al. (2024). Accuracy is evaluated with task-oriented, reference-based metrics including ROUGE-2, ROUGE-L, F1 He et al. (2024), and factuality metric AlignScore Zha et al. (2023), using natural text as the ground truth.
In Step 2, we evaluate the overall quality by proposing a quality metric Factuality-weighted Score to see how the watermarking method affects factuality.
Factuality-weighted Score (FWS)  The proposed aspects allow us to quantify the overall quality metric as shown in Equation 1. To prioritize relevant context and factual accuracy over coherence, we set $\alpha=0.4$ and $\beta=0.2$ to indicate that Relevance or Completeness ($Rel.$) and Accuracy ($FactAcc.$) are equally important, and they are twice as important as Coherence ($Coh.$). This weighting was based on empirical consideration through sensitivity analysis provided in Table 5.
$FWS=\alpha(\text{Rel}+\text{FactAcc})+\beta(\text{Coh})$ (1)
@@ 3 Evaluation Workflow
Human evaluation  To evaluate both automatic metrics and GPT-Judger ratings on equal footing, we conducted a human evaluation with 6 respondents, including graduate students and medical practitioners, ensuring familiarity with LLMs and medical expertise. The evaluation focused on the KGW watermarking method applied to the Question Answering (QA) task, chosen as KGW underlies many LLM watermarking approaches and QA represents the most realistic practical use case for medical language models. Each item was rated by 3 respondents to obtain an average score using the quality aspects in Table 1. The full evaluation template is provided in Appendix B.2.
Pearson correlation was computed to measure agreement between human judgments and both traditional automatic metrics and GPT-Judger ratings. Prior to this, a Nemenyi test was applied to detect significant differences among the proposed aspects, with lower p-values indicating greater distinction Fu et al. (2024).
@@ 4.2 Watermarking Methods
We evaluate six watermarking methods across two categories Liu et al. (2024). Post-hoc watermarking applies after text generation. We evaluate two schemes from this category: LogRank Gehrmann et al. (2019), which classifies text by thresholding the average log-rank of tokens, as machine-generated text typically has smaller average log-ranks, and DetectGPT Mitchell et al. (2023), which exploits the tendency of model-generated text to occupy negative curvature regions in the log probability function by comparing original passages with semantically similar perturbations.
The second category is generation-time watermarking, where watermarks are embedded during the LLM’s text generation process. We evaluate four approaches: KGW Kirchenbauer et al. (2023), a foundational logit-based method that partitions the vocabulary into red and green lists based on a previous token hash; SWEET Lee et al. (2024), an improvement of the KGW method on addressing low entropy tokens for code generation domain However, another work Huo et al. (2024) showed that SWEET works well in the general domain, thus we believe potentially useful in medical text where low entropy tokens can be found; DiPmark Wu et al. (2024) a distribution-preserving watermark that uses complementary reweighting strategies with random token permutations to maintain original text quality; and EXP-edit Kuditipudi et al. (2024), a token sampling-based method modifying probabilities based on token rank.
This selection allows comparison across three strategies post-hoc (PH), logit-based (LB), distribution preserving or distortion-free (DF), while also includes methods addressing potential domain-specific challenges like low entropy. All generation-time watermarks used the MarkLLM toolkit Pan et al. (2024) with default parameters. Brief methods are available in Appendix B.4, and parameter details are available in Appendix A.2, Table 10.
@@
4.3 Tasks and Datasets
To assess watermark performance in realistic medical scenarios, we employ three distinct tasks using established medical datasets. First, Text Completion using the HealthQA dataset Zhu et al. (2019) requires models to complete medical passages (approx. 200 tokens), evaluating the impact on fundamental coherence and fluency. Second, Question Answering (QA), also on HealthQA, challenges models to answer medical questions (approx. 200 tokens), specifically testing factual accuracy preservation under watermarking. Third, Summarization using the MeQSum dataset Abacha and Demner-Fushman (2019) involves generating concise summaries (< 100 tokens) of medical questions, evaluating performance on challenging short-text generation Kirchenbauer et al. (2023) and the crucial preservation of essential factual information. These tasks represent diverse medical tasks and output lengths, offering a comprehensive view of watermark impact on text quality and factual integrity in the medical domain. Details on how we built each task are available in Appendix B.3.
4.4 Model and Hardware
We use Meditron-7B Chen et al. (2023) as the main model for most experiments. This model included training on clinical guidelines, more suitable for consumer QA, unlike most models, which limited their training on abstracts and articles. Meditron is a medical adaptation model from Llama-2-7B Touvron et al. (2023) through continued pretraining on the medical GAP-Replay corpus. We also evaluate other medical models like MedLlama-3-8B [^3] https://hf.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0 and BioMistral 7B Labrak et al. (2024). Though this variation is not our main focus, we only include RQ1 results using these models and provide additional analysis in Appendix A.1. All experiments are conducted on a workstation equipped with an AMD Ryzen Threadripper PRO 3955WX (16-Cores) processor and two NVIDIA GeForce RTX 3090 GPUs (24GB VRAM each).
5 Results
@@ 5.1 Detectability and Quality (RQ1)
Table 2 presents comparative performance of watermarking methods across generation tasks, measuring both detection capability and text quality preservation. TPR measures detection rate, AUROC overall robustness, PPL text fluency, and SimCSE semantic similarity. The results show a clear trade-off between detectability and output quality among different watermarking approaches.
@@ 5.2 Task-oriented Performance (RQ2)
Table 3 presents reference-based quality metrics evaluating generated text against natural text ground-truth, deliberately to give insights of factual preservation across different tasks. Unlike the general quality metrics in RQ1, task-specific metrics provide finer granularity ROUGE-2 measures bigram overlap between generated and reference text, ROUGE-L captures the longest common subsequence to reflect order matching, F1 balances precision and recall to quantify overall matching accuracy and AlignScore evaluates factual consistency. Higher values across all metrics indicate better performance.
@@ 5.3 Factuality-weighted Score FWS (RQ3)
For the task-specific performance across watermarking approaches: SWEET demonstrates superior performance in Text Completion (0.556), while KGW achieves the highest scores in both Question Answering (0.411) and Summarization (0.419). EXP-edit consistently shows the weakest performance across all tasks. This is likely due to its alternative token sampling strategy, which causes the model to behave almost like a different language model compared to the un-watermarked version. In this case, it introduces more drawbacks than the intended benefit of distortion-free watermark, as claimed Kuditipudi et al. (2024).
@@
Configuration
FWS Param.
($\alpha$, $\beta$)
Auto-
Human
GPTJ-
Human
Coherence-Heavy (1:2)
(0.25, 0.5)
0.089 0.814
Equal Weighting (1:1)
(0.33, 0.33)
0.181 0.833
Current (Ours) (2:1)
(0.4, 0.2)
0.248 0.840
Factuality-Heavy (4:1)
(0.44, 0.11)
0.276 0.841
Factuality-Heavy (6:1)
(0.46, 0.08)
0.283 0.841
@@
Human Evaluation  Human ratings were evaluated using the Nemenyi test and p-value significance on quality aspects defined in Table 1. The p-values for (Coherence, Relevance), (Coherence, FactualAccuracy), and (Relevance, FactualAccuracy) are 0.635, 0.001, and 0.004, indicating clear distinctions among the aspects. These quality aspects exhibit distinct characteristics consistent with intuition, highlighting the importance of detailed factuality aspects in the medical domain. Sensitivity analysis of Equation 1 shows a consistent trend: greater emphasis on factuality improves correlation with human judgment. While the 2:1 weighting did not yield the highest correlation, it outperformed coherence-heavy settings and was chosen as an intuitive benchmark for a factuality-focused configuration (Table 5).